# Accelerating Speculative Decoding using Dynamic Speculation Length

### v1

@@ Abstract
Speculative decoding is a promising method for reducing the inference latency of large language models. The effectiveness of the method depends on the speculation length (SL)—the number of tokens generated by the draft model at each iteration. The vast majority of speculative decoding approaches use the same SL for all iterations. In this work, we show that this practice is suboptimal. We introduce DISCO, a DynamIc SpeCulation length Optimization method that uses a classifier to dynamically adjust the SL at each iteration, while provably preserving the decoding quality. Experiments with four benchmarks demonstrate average speedup gains of 10.3% relative to our best baselines. [^1] Code will be available upon publication.
Accelerating Speculative Decoding using Dynamic Speculation Length
@@ 1 Introduction
Large language models (LLMs) generate tokens in an autoregressive manner, incurring slow and costly generation. Speculative decoding methods Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023) mitigate this issue by splitting the inference of a given target model into two steps. In the first step, a faster draft model generates tokens autoregressively. In the second step, the target model validates all the draft tokens simultaneously. The validation procedure accepts or rejects the draft tokens (Fig. 1). This approach accelerates the generation process by producing multiple tokens in a single forward pass of the target model, while maintaining the decoding quality of the target model.
The effectiveness of speculative decoding in expediting the token generation process depends on the speculation length (SL) parameter, which pertains to the number of tokens generated by the draft model before each validation step (Chen et al., 2023). The choice of SL has a large effect on latency: lower static (constant) SL values may incur a lower number of accepted draft tokens per speculative iteration, which leads to more target model validations than necessary. In contrast, higher static SL may result in a large number of ignored draft generations in case of an early draft-target mismatch (Fig. 1).
Figure 1: An illustration of a single speculative decoding iteration with static speculation length ($SL$) $=5$. A draft model generates $SL$ tokens autoregressively. The target model validates them all in parallel, and accepts the first three tokens. In this case, $SL$ is too high (as the last two draft tokens are never used). $SL$ can also be too low (e.g., $SL=2$ here), which would lead to too many calls to the target model.
The vast majority of speculative decoding approaches Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023); Cai et al. (2024); Li et al. (2024) employ a static SL, i.e., a constant number of draft tokens are generated in each speculative iteration. In this work, we show that this approach is suboptimal. We start by introducing a dynamic oracle SL, which sets the SL to be the exact number of draft generations until the first mismatch in each speculative iteration. We observe a high variance in the optimal SL across the speculative iterations. This finding demonstrates the potential of dynamically adapting the speculation length.
We then propose DISCO, a dynamic speculation length optimization method. DISCO uses draft model features to train a lightweight classifier that predicts the compatibility between a generated draft token and the target token. At inference time, the classifier determines whether the draft model should continue generating the next token, or halt and switch to the target model for validation.
We evaluate DISCO’s effectiveness in code generation, text summarization and instruction tasks. Our results demonstrate average speedup gains of 10.3% and 31.4% relatively to the best static and dynamic SL baselines, respectively, while maintaining the same generated text. Our classifier also transfers well across tasks from the same category: training it on one task and using it on another leads to similar gains over our baselines.
Figure 2: The improvement factor (IF) as a function of the speculation length ($\gamma$) for three different acceptance rate ($\alpha$) values assuming fixed cost coefficient ($c=0.05$). The IF value largely depends on both $\alpha$ and $\gamma$.
@@ 2 Background: Speculative Decoding
Speculative decoding Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023) expedites LLM generation while ensuring no accuracy loss by dividing it into two stages. In the first stage, a fast but less accurate draft model $M_{D}$ autoregressively generates a sequence of tokens. In the second stage, a large but more accurate target model $M_{T}$ conducts parallelized verification over the generated draft tokens. This process allows the model to potentially produce multiple tokens per target forward pass.
Speculation length (SL)
The effectiveness of speculative decoding in accelerating the token generation process relies heavily on the SL parameter, which determines how many tokens are generated by the draft model before each validation step. The effect of the SL on the overall speedup is subject to a tradeoff; higher SLs potentially reduce the number of target model validations, but also increase the number of redundant draft generations (Fig. 1).
The majority of speculative decoding approaches Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023); Cai et al. (2024); Li et al. (2024) use a static SL—the same number of draft tokens are generated per speculative iteration.222 Some practical work apply heuristics for dynamic SL adaptation, increasing/decreasing SL when all/some draft tokens are accepted/rejected in the previous iteration (https://huggingface.co/blog/assisted-generation). Chen et al. (2023) explored various static SLs, across different target-draft model pairs and tasks, and empirically showed that as the SL rises, the overall speedup increases until reaching a certain threshold, beyond which it either levels off or even regresses.
To study the effect of different static SL values, Leviathan et al. (2023) defined the improvement factor (IF) as the expected token generation speedup. Formally:
@@
where $\alpha$ denotes the acceptance rate, indicating the expected probability of a draft token to be accepted by the target model; $c$ represents the cost coefficient, indicating the ratio between the walltime of a forward pass run of the draft model $M_{D}$ and the wall time of a forward pass run of the target model $M_{T}$; and $\gamma$ represents the static SL value. [^3] IF computation assumes enough compute resources for increased concurrency as $\gamma$ rises.
Figure 2 shows the IF as a function of $\gamma$ for three different values of $\alpha$ and a fixed $c$ value. While both $\alpha$ and $c$ are important to the selection of the target-draft model pair, finding the optimal $\gamma$ is fundamental to the effectiveness of the system.
3 Dynamic Speculation Length
@@
The IF function is based on the simplifying assumption that the probability of accepting draft tokens by the target model is i.i.d. Nevertheless, in practical scenarios, different tokens may have varying levels of predictability, rendering the simplifying i.i.d. assumption unsound. Therefore, relying on a static SL might not be optimal, as certain speculative iterations may benefit from generating more tokens than others. Below we consider an oracle experiment, which applies the optimal dynamic $\gamma$ value at each iteration. We then propose a method for dynamically setting $\gamma$, showing that it strongly outperforms a static selection method for any choice of a static SL.
Assessing the optimal dynamic $\gamma$ value
@@
We start by employing an oracle for detecting the ‘ground truth’ value of $\gamma$ for each speculative iteration. The oracle uses the draft model to autoregressively generate tokens until a mismatch occurs between the predicted tokens of the draft and target models. This process is repeated for each speculative iteration, ultimately returning the optimal (maximum) number of accepted draft tokens per iteration. The mismatch between the tokens is determined by using the rejection sampling algorithm introduced by Leviathan et al. (2023) with zero temperature. This oracle fulfills the speculative decoding potential: generating the maximal number of valid draft tokens at each iteration, and making a minimal number of calls to both draft and target model.
@@
Figure 3 shows the oracle SL values across the speculative iterations for one MBPP example. Comparing with static SL, we observe a lower number of oracle draft and target forward passes. Figure 4 shows the average oracle SL over the speculative iterations for Alpaca dataset. We observe on both figures a high variance of oracle SL values, implying that setting a predefined static SL is likely to be suboptimal. See App. A for further analysis.
@@
Figure 4: Bar chart of the average oracle SL over the normalized index of the speculative iterations for the Alpaca dataset. We observe a high variance of oracle SL values.
DynamIc SpeCulation length Optimization
@@
We introduce DISCO, a simple method for dynamically setting the SL value at each iteration. To estimate the correct SL value at each step, we employ a simple SL classifier. Our classifier decides, at each draft generation step, whether the draft model $M_{D}$ should proceed to generate the next token or switch to the target model $M_{T}$ for verification. The classifier takes as inputs the probability vector of the draft model ($y^{D}_{i}$) and the token position ($i$), generating a confidence score ($C_{i}$) used for the decision-making as follows:
@@
where $Top_{k}()$ selects the top $k$ values [^4] We use $k=10$ in all experiments. and $Ent()$ is the entropy function.
@@ 4 Experiments
We evaluate our method on four datasets spanning three tasks: code generation, using MBPP Austin et al. (2021) and HumanEval Chen et al. (2021); text summarization using CNN-DailyMail Nallapati et al. (2016); and instruction-following using Alpaca Taori et al. (2023). We use the training set for training the SL classifier and the validation set for setting threshold $\tau$ and $SL_{max}$ hyperparameters. For HumanEval, which has no training and validation sets, we evaluate transfer learning from MBPP. See App. B for more details on the datasets and prompts.
@@
To evaluate speculative decoding on code generation tasks, we use the Starcoder model family Li et al. (2023)—15B for target and 168M for draft. For our text generation tasks, we use Vicuna models—13B as target Chiang et al. (2023) and 68M as draft Yang et al. (2024). Throughout the experiments we use the rejection sampling scheme Leviathan et al. (2023) with the temperature set to zero. Exact model names and versions are provided in App. C.
@@
For minimal overhead, we use a shallow 2-layer FFN classifier. We train the SL classifier to predict the conformity between $M_{D}$ and $M_{T}$ by employing cross-entropy loss with the Total Variance distance: $TV(y^{D}_{i},y^{T}_{i})$, where $y^{D}_{i}$ and $y^{T}_{i}$ represent the vocabulary distribution of $M_{D}$ and $M_{T}$ respectively at the $i^{th}$ token position. We evaluate the quality of the classifier by measuring its F1 score on the validation set. The F1 results obtained on the datasets are relatively high; for instance, the F1 score on the MBPP dataset is 95% while the F1 score of the optimal static SL is 85%. For further details, see App. E and App. D.
@@
We compare the LLM inference latency of DISCO to both static SL (Static SL-5) and dynamic heuristic SL setups (dynHeur SL; see Fn. 2). We also consider the optimal static SL baseline tuned on our validation sets (Static SL-opt). Finally, we compare DISCO to our oracle (Sec. 2) that represents the lower bound on latency.
@@ Results
Dataset Method Latency(ms) Speedup
@@
HumanEval* Target 23.46 1.00x
@@
Table 1: Average latency results on different benchmarks. HumanEval results (marked with *) use a classifier trained on MBPP (transfer learning).
Table 1 presents our results. Employing an SL classifier consistently outperforms all other baselines across all benchmarks. Average latency improvements of DISCO over the optimal static SL and the dynamic heuristic baselines are 10.3% and 31.4% respectively (see App. F), while preserving the same output as the target model. Importantly, our improvement does not come only from our training data: the optimal static SL (as fit by that data) is still underperformed by DISCO. Finally, we note that classifiers transfer well across tasks: when trained on MBPP, DISCO is still able to outperform all baselines on HumanEval.
@@ 5 Related Work
There has been a growing body of work in the emerging field of speculative decoding for LLM acceleration. Pioneering studies by Leviathan et al. (2023) and Chen et al. (2023) introduced a novel rejection sampling scheme which preserves the distribution of the target model, effectively guaranteeing that speculative decoding setups will maintain the quality of the target model. Subsequent work Miao et al. (2024) elevated the average number of accepted tokens by using several draft models. More recent works Cai et al. (2024); Li et al. (2024); Bhendawade et al. (2024) eliminated the need for a separate draft model by incorporating specialized draft layers on top of the transformer decoder to forecast a sequence of draft tokens. All of the above works employed a static SL.
A very recent concurrent work by S et al. (2024) enhanced the draft model’s accuracy by granting it access to the target model’s representations. In addition, it employed a classifier to determine whether to halt or continue the speculation process. In this sense, this work is the closest to ours; however, our work delves into the impact of the SL on the efficiency of speculative decoding, encompassing comparisons between static and dynamic SL approaches, as well as the upper bound of improvement represented by the oracle SL.
@@ 6 Conclusion
We introduced DISCO, a dynamic speculation length optimization method. The method uses a classifier that determines whether the draft model should continue to generate the next token or pause and transition to the target model for validation. We evaluated DISCO’s effectiveness using four benchmarks and demonstrated average speedup gains of 10.3% and 31.4% relatively to the optimal static SL and dynamic heuristic baselines. Our results highlight the potential of further reducing inference cost by using simple, efficient techniques.
@@ 7 Limitations
Second, we note that the effectiveness of dynamic speculation length (SL) approaches is contingent upon the standard deviation of the SL as measured by the oracle. When the standard deviation is low, there may be less room for improvement using dynamic SL strategies.

### v4

# Dynamic Speculation Lookahead Accelerates Speculative Decoding of Large Language Models

@@ Abstract
Speculative decoding is commonly used for reducing the inference latency of large language models. Its effectiveness depends highly on the speculation lookahead (SL)—the number of tokens generated by the draft model at each iteration. In this work we show that the common practice of using the same SL for all iterations (static SL) is suboptimal. We introduce DISCO (DynamIc SpeCulation lookahead Optimization), a novel method for dynamically selecting the SL. Our experiments with four datasets show that DISCO reaches an average speedup of 10% compared to the best static SL baseline, while generating the exact same text. [^1] We will publicly release our code upon publication.
Dynamic Speculation Lookahead Accelerates
Speculative Decoding of Large Language Models
@@ 1 Introduction
Large language models (LLMs) generate tokens autoregressively (Radford et al., 2019; Brown et al., 2020), which often leads to slow generation. Speculative decoding algorithms Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023) reduces the inference latency by splitting the inference of a given model into two steps. First, a fast draft model generates tokens autoregressively. Then a more accurate (target) model validates all generated draft tokens simultaneously. See Fig. 1 for an example. The effectiveness of speculative decoding depends on the speculation lookahead (SL)—the number of tokens generated by the draft model at each iteration. An SL too small leads to too many target forwards steps; SL values too large add redundant draft forward passes. Yet, existing speculative decoding approaches use a static SL—an SL that remains constant across all iterations (Leviathan et al., 2023; Miao et al., 2024; Chen et al., 2023; Cai et al., 2024; Li et al., 2024).
Figure 1: An illustration of a single speculative decoding iteration with Speculation Lookahead (SL) = 5. Given a prompt $t_{0}$, a draft model autoregressively generates 5 tokens $t_{1},\ldots,t_{5}$. The target model validates them all in parallel and accepts only $t_{1},t_{2},t_{3}$. As $t_{4}$ and $t_{5}$ are rejected, the SL is suboptimal (too large).
This work starts by defining an oracle SL—a method that assigns each iteration its optimal SL. We observe that the optimal SL shows a high variance across iterations (Fig. 2). We then use this oracle to estimate an upper bound of the expected speedup (compared to using static SLs), showing a potential gain of up to 39% speedup. We then propose DISCO, a novel method for selecting the SL before each iteration. DISCO estimates the likelihood of the next draft token being accepted by the target model, and halts the draft model if this likelihood is too small. We evaluate DISCO across various tasks: code generation, text summarization, and instruction following. Our results show an average speedup of 10% compared to optimal static SL and 31% compared to a previously known heuristic for controlling SLs Gante (2023), all without modifying the output text (Appendix F). DISCO also transfers well across tasks from the same category: training it on one task and using it on another leads to similar speedups.
@@ 2 Background: Speculative Decoding
Speculative decoding expedites LLM generation while ensuring no accuracy loss by dividing it into two stages. In the first stage, a fast but less accurate draft model $M_{D}$ autoregressively generates a sequence of tokens. In the second stage, a large but more accurate target model $M_{T}$ conducts parallelized verification over the generated draft tokens. This process allows the model to potentially produce multiple tokens per target forward pass.
Speculation lookahead (SL)
The effectiveness of speculative decoding in accelerating the token generation process relies heavily on the SL parameter, which determines how many tokens are generated by the draft model before each validation step. The effect of the SL on the overall speedup is subject to a tradeoff; higher SLs potentially reduce the number of target model validations, but also increase the number of redundant draft generations (Fig. 1), and vice-versa. The majority of speculative decoding approaches Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023); Cai et al. (2024); Li et al. (2024) use a static SL—the same number of draft tokens are generated per speculative iteration. [^2] A notable exception is Gante (2023), who applies a heuristic for dynamic SL adaptation by modifying the SL based on the acceptance rate of previous iterations. Chen et al. (2023) explored various static SLs, across different target-draft model pairs and tasks, and empirically showed that as the SL rises, the overall speedup increases until reaching a certain threshold, beyond which it either levels off or even regresses. To study the effect of the SL, Leviathan et al. (2023) defined the improvement factor (IF) as the expected token generation speedup:
@@
where $\alpha$ denotes the acceptance rate, indicating the expected probability of a draft token to be accepted by the target model; $c$ represents the cost coefficient, indicating the ratio between the walltime of a forward pass run of the draft model $M_{D}$ and the wall time of a forward pass run of the target model $M_{T}$; and $\gamma$ represents the static SL value. [^3] IF computation assumes enough compute resources for increased concurrency as $\gamma$ rises. While both $\alpha$ and $c$ are important to the selection of the target-draft model pair, finding the optimal $\gamma$ is fundamental to the effectiveness of the system.
3 Dynamic Speculation Lookahead
@@
The IF function (Eq. 1) is based on the simplifying assumption that the probability of accepting draft tokens by the target model is i.i.d. Nevertheless, in practical scenarios, different tokens may have varying levels of predictability, which challenges this i.i.d. assumption, and suggests that using a static SL might be suboptimal. Below we consider an oracle experiment, which applies the optimal dynamic $\gamma$ value at each iteration. We then propose a method for dynamically setting $\gamma$, showing that it strongly outperforms a static selection method for any choice of a static SL.
Finding the optimal SL per iteration
@@
We start by employing an oracle for detecting the optimal value of SL ($\gamma$) for each speculative iteration. The oracle uses the draft model to autoregressively generate tokens until a mismatch occurs between the predicted tokens of the draft and target models. This process is repeated for each speculative iteration, ultimately returning the optimal (maximum) number of accepted draft tokens per iteration. The mismatch between the tokens is determined by using the rejection sampling algorithm introduced by Leviathan et al. (2023) with zero temperature. This oracle fulfills the speculative decoding potential: generating the maximal number of valid draft tokens at each iteration, and making a minimal number of calls to both draft and target model.
@@
Figure 2 shows the oracle SL values across the speculative iterations for one MBPP example. Compared to the static SL, we observe a lower number of both draft and target forward passes. Figure 3 shows the average oracle SL over the speculative iterations for the Alpaca dataset Taori et al. (2023). Both figures show a high variance of oracle SL values, implying that astatic SL is likely to be suboptimal. See App. A for further analysis.
@@
Figure 3: The average oracle SL over the normalized index of the speculative iterations for the Alpaca dataset. We observe a high variance of oracle SL values.
DynamIc SpeCulation lookahead Optimization
@@
We introduce DISCO, a simple method for dynamically setting the SL value at each iteration. To estimate the correct SL value at each step, we employ a simple classifier as follows. Immediately after generating any draft token, our classifier decides whether the draft model $M_{D}$ should proceed and generate the next token or switch to the target model $M_{T}$ for verification. The classifier takes as inputs the probability vector of the draft model ($y^{D}_{i}$) and the token position ($i$), generating a confidence score ($C_{i}$) used for the decision-making as follows:
@@
where $Top_{k}()$ selects the top $k$ values and $Ent()$ is the entropy function. [^4] We use $k=10$ in all experiments.
@@ 4 Experiments
We evaluate our method on four datasets spanning three tasks: code generation using MBPP Austin et al. (2021) and HumanEval Chen et al. (2021); text summarization using CNN-DailyMail Nallapati et al. (2016); and instruction-following using Alpaca. We use the training sets for training the SL classifier and the validation sets for setting the threshold $\tau$ and the $\text{SL}_{\max}$ hyperparameter. For HumanEval, which has no training and validation sets, we evaluate transfer learning from MBPP. See App. B for more details on the datasets and prompts.
@@
For code generation tasks, we use the Starcoder model family Li et al. (2023)—15B for target and 168M for draft. For the other tasks, we use Vicuna models—13B as target Chiang et al. (2023) and 68M as draft Yang et al. (2024). See App. C for more details.
@@
To train the classifier we extract features from the training sets of our datasets MBPP, CNN-DM, and Alpaca. For minimal overhead, we use a shallow 2-layer FFN classifier and train it based on the extracted features to predict the agreement between the draft and target models. The training employs cross-entropy loss with Total Variance as distance measure: $TV(y^{D}_{i},y^{T}_{i})$, where $y^{D}_{i}$ and $y^{T}_{i}$ represent the vocabulary distribution of $M_{D}$ and $M_{T}$ respectively at the $i^{th}$ token position. We evaluate the quality of the classifier by measuring its F1 score on the validation set. The F1 results obtained on the datasets are relatively high; for instance, 95% on MBPP, compared to 85% using the optimal static SL. See Apps. D and E for more details.
@@
We compare the LLM inference latency of DISCO to both static SL (static SL-5) and dynamic heuristic SL setups (dynHeur SL; Gante, 2023). We also consider the optimal static SL baseline tuned on our validation sets (static SL-opt). Finally, we also report results for our oracle (Sec. 2), which represents the lower bound on latency.
@@ Results
Benchmark Method Latency(ms) Speedup
@@
HumanEval (transfer learning) Target 23.46 1.00x
@@
Table 1: Average latency results on different benchmarks. HumanEval results use a classifier trained on MBPP (transfer learning). All results are provided with greedy decoding (temperature=0).
Table 1 presents our results using the rejection sampling scheme with greedy decoding (temperature=0) since baselines get higher speedup Leviathan et al. (2023). Employing an SL classifier consistently outperforms all other baselines across all benchmarks. Average latency improvements of DISCO over the optimal static SL and the dynamic heuristic baselines are 10.3% and 31.4% respectively, while preserving the same output as the target model. Importantly, our improvement does not come only from our training data: the optimal static SL (as fit by that data) is still underperformed by DISCO. Finally, DISCO transfers well across tasks: when trained on MBPP, it is still outperforms all baselines on HumanEval.
Analysis
Since DISCO is sampling temperature agnostic, we present in Table 2 latency results for CNN-DM with non-zero temperature; we observe that DISCO speedup also improves in that case. In addition, we report latency results using a simple rule-based approach: we use the perplexity to measure the confidence in the sequence of tokens predicted by the draft, when lower perplexity indicates higher confidence. Optimal perplexity threshold is tuned on our validation sets (ppl SL-opt). We observe that it yields a lower speedup.
Benchmark Method Latency(ms) Speedup
CNN-DM Target 36.32 1.00x
static SL-5 21.88 1.66x
DISCO (ours) 19.96 1.82x
ppl SL-opt 26.79 1.43x
Table 2: Additional average latency results for CNN-DM: temperature=1; perplexity-based baseline.
@@ 5 Related Work
Pioneering studies on speculative decoding Leviathan et al. (2023); Chen et al. (2023) introduced a rejection sampling scheme that preserves the distribution of the target model, guaranteeing that speculative decoding maintains the quality of the target model. Subsequent work Miao et al. (2024) elevated the average number of accepted tokens by using several draft models. Most recently, Timor et al. (2024) introduced DSI, a distributed variation of speculative decoding that is provably faster than non-distributed methods and does not require additional training or architectural changes. DSI orchestrates multiple servers, and the maximum number of servers is inversely proportional to the SL. To eliminate the need for a separate draft model, Li et al. (2024); Cai et al. (2024); Bhendawade et al. (2024); Yang et al. (2024) train additional, specialized draft layers on top of the transformer decoder. DISCO transfers well within domains and does not require classifier training per dataset whereas these methods necessitate training per dataset. At inference time, they employ a static SL; we believe that DISCO can be beneficially applied to these approaches, we leave this research for future work. Zhang et al. (2023) proposed draft-exiting with an adaptive threshold for self-speculative decoding using a rule-based approach that compares a confidence to a predetermined threshold. This method seems suitable to approaches where the draft is a subset of the target model, whereas our approach is more generic. A very recent concurrent work by S et al. (2024) enhanced the draft model’s accuracy by granting it access to the target model’s representations. In addition, it employed a classifier to determine whether to halt or continue the speculation process. Our work delves into the impact of the SL on the efficiency of speculative decoding, encompassing comparisons between static and dynamic SL approaches, as well as the upper bound of improvement represented by the oracle SL.
@@ 6 Conclusion
We have shown that using the same speculation lookahead parameter across speculative decoding iterations is suboptimal. We introduced DISCO, a dynamic speculation lookahead optimization method. The method uses a classifier that determines whether the draft model should continue to generate the next token or pause and transition to the target model for validation. We evaluated DISCO’s effectiveness using four benchmarks and demonstrated average speedup gains of 10.3% and 31.4% relatively to the optimal static SL and dynamic heuristic baselines. Our results highlight the potential of further reducing inference cost by using simple, efficient techniques.
@@ 7 Limitations
Second, we note that the effectiveness of dynamic speculation lookahead (SL) approaches is contingent upon the standard deviation of the SL as measured by the oracle. When the standard deviation is low, there may be less room for improvement using dynamic SL strategies.